# Pcodec: Better Compression for Numerical Sequences

### v1

@@ 3.2.1 Setup
All benchmarks were performed on a bare metal Intel Xeon Platinum 8488C CPU. We ran both single- and multi-threaded benchmarks. Our multi-threaded benchmarks were designed to simulate moderate load, using 48 of the 96 cores such that each active core had its own L2 cache. To minimize measurement noise, we took the median of at least 3 iterations after a warm-up iteration. We limited large datasets to 2 million rows.
We compared Pco’s default configuration versus Blosc2, Parquet, SPDP, Turbo PFor, and Zstd. For Blosc2 we used Shuffle, and for Parquet we considered both Delta and Dictionary. Since Parquet Delta only applies to integers, we always used Dictionary for floats. Blosc2, Parquet, and Turbo PFor were paired with Zstd. We also tested LZ4 and Brotli codecs in place of Zstd, but found that Zstd was strictly better on these datasets.
@@ 48 Zstd 1,178 839 768 893 1,132 912
Table 3: Single-threaded and multi-threaded decompression speeds per thread in MiB/s. Decompression speed varies only slightly between compression levels, so we report its harmonic mean over compression levels. Pco’s consistently decompressed over 1GiB/s, even in multi-threaded settings.
@@ 3.2.2 Interpretation
Given similar compression time, we found that Pco compressed substantially better than all alternatives on all datasets. Specifically, at default compression level, Pco compressed 29% to 94% better while using less compression time (Table 4). This matched our intuition that binning would suit numerical data better than LZ codecs do.
@@ 4 Conclusion
Pcodec is a mature compression format that offers better compression than alternatives. It accomplishes this via a novel binning procedure, distinct from LZ-based approaches.
@@ Appendix A Derivation of Binning Bit Cost Bound
Theorem 1 Suppose $X$ is mixture over $\{0,\ldots,T-1\}$ of $s$ disjoint integer distributions, each of which has a monotonic PMF. Then for any $k>2s$, there exists a binning of at most $k$ bins such that the binned bit cost $\hat{H}$ of a random draw from $X$ satisfies
$\hat{H}\leq H+\frac{5s\log_{2}(T)}{k-2s}$
@@
And less formally, we will show that this binning roughly consists of bins of equal probability. Also informally, some bounds in this appendix were made looser than necessary for the sake of simplicity. The constant $5$ could be brought down to $4+\epsilon$ by imposing a lower bound on $T$, and the scaling behavior with imbalanced mixture distributions could both be improved in subsection A.4.
To sketch the proof, there will be three main steps. First, we will show that there exists a binning in which every nontrivial bin has probability at most $2/(k-1)$. Next, we will bound the bit cost of a monotonic distribution using such a binning. Finally, we will extend this to mixtures of monotonic distributions.
@@ A.1 Definitions
Let $P(x)$ be the PMF of a random integer variable $X$ over a finite, contiguous domain of $T$ distinct $x$ values. Additionally, let a binning with $n$ bins consist of a list of intervals $([a_{1},b_{1}],\ldots,[a_{k},b_{k}])$. Each such bin has various properties:
@@ A.2 Constructive Bound on Bin Weight
For any $k$, there exists a binning of $X$ with $k$ or fewer bins such that each bin is either trivial or has weight at most $2/(k-1)$.
We will prove this constructively. First, for any $x$ such that $P(x)>2/(k-1)$, dedicate a trivial bin to that single $x$ value. There can be at most $(k-1)/2$ of these, so suppose there are $k_{\text{trivial}}$. In between these, there can be at most $k_{\text{trivial}}+1$ nontrivial intervals of numbers that do not yet have bins assigned.
Consider one of these nontrivial intervals. We assign bins over it greedily, until their weight would exceed $2/(k-1)$:
@@
  • For each $x$ value, let $w_{\text{proposed}}=w_{\text{current}}+P(x)$, and
    • If $w_{\text{proposed}}\leq 2/(k-1)$, assign $w_{\text{current}}\leftarrow w_{\text{proposed}}$ and add $x$ to the current bin.
@@
Clearly each such bin has weight mass of at most $2/(k-1)$. And every pair of such bins must have weight over $2/(k-1)$ or else they would have been combined during the greedy algorithm. Thus if there are $k_{\text{nontrivial}}$ consecutive nontrivial bins, their weight must be at least $(2/(k-1))\lfloor k_{\text{nontrivial}}/2\rfloor$, which is at least $(2/(k-1))\cdot(k_{\text{nontrivial}}-1)/2$ since $k_{\text{nontrivial}}$ is a positive integer.
The weight of the trivial bins is at least $(2/(k-1))k_{\text{trivial}}$, and the weight of the nontrivial bins is at least $\sum_{i=0}^{k_{\text{trivial}}}(2/(k-1))\cdot(k_{\text{nontrivial},i}-1)/2$. Since the total probability is 1, we must have
$\frac{2}{k-1}\left(k_{\text{trivial}}+\sum_{i=0}^{k_{\text{trivial}}}(k_{\text{nontrivial},i}-1)/2\right)\leq 1$
$\frac{2}{k-1}\left(\frac{1}{2}k_{\text{trivial}}-\frac{1}{2}+\sum_{i=0}^{k_{\text{trivial}}}(k_{\text{nontrivial},i})/2\right)\leq 1$
$k_{\text{trivial}}+\sum_{i=0}^{k_{\text{trivial}}}k_{\text{nontrivial},i}\leq k$
In other words, we have used at most $k$ bins, each of which is either trivial or used weight at most $2/(k-1)$. This completes the proof of Lemma 1.
@@ A.3 Bound for Monotonic PMFs
$\hat{H}\leq H+\frac{5\ln(T)}{k-1}.$
We will show that the binning from Lemma 1 satisfies this.
Without loss of generality, assume $X$’s PMF is monotonically increasing. Consider a bin $[a_{i},b_{i}]$. The partial entropy of this bin is
@@
Now we need only consider nontrivial bins. We will further divide these into shallow and deep bins. Let $k_{\text{shallow}}$ be the smallest value such that $\sum_{i=1}^{k_{\text{shallow}}}w_{i}\geq\frac{1}{k-1}$ (if no such value exists, then we have essentially only trivial bins, and it is easy to show that the lemma holds). We can rewrite the inequality as
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\sum_{i=k_{\text{shallow}}+1}^{k_{\text{nontrivial}}}w_{i}\left(\log_{2}(T_{i})-\log_{2}(w_{i})+\log_{2}(q_{i})\right)$ (3)
@@
We can use this to bound the deep portion of 3,
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\sum_{i=k_{\text{shallow}}+1}^{k_{\text{nontrivial}}}w_{i}\left(\log_{2}\left(\frac{w_{i}}{q_{i-1}}\right)-\log_{2}(w_{i})+\log_{2}\left(q_{i}\right)\right)$
@@
Since each of these bins has $w_{i}\leq\frac{2}{k-1}$,
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\frac{2}{k-1}\sum_{i=k_{\text{shallow}}+1}^{k_{\text{nontrivial}}}\log_{2}\left(q_{i}\right)-\log_{2}(q_{i-1})$
@@
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\frac{2}{k-1}\left(\log_{2}\left(q_{k_{\text{nontrivial}}}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
Since $q_{i}\leq\frac{2}{k-1}$ for deep $i$,
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\frac{2}{k-1}\left(\log_{2}\left(\frac{2}{k-1}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
@@
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}w_{i}\log_{2}(T_{i})+\frac{2}{k-1}\left(\log_{2}\left(\frac{2}{k-1}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
@@
$\hat{H}-H\leq w_{\text{shallow}}\log_{2}(T)+\frac{2}{k-1}\left(\log_{2}\left(\frac{2}{k-1}\right)-\log_{2}\left(q_{k-\text{shallow}}\right)\right)$
@@
$\hat{H}-H\leq w_{\text{shallow}}\log_{2}(T)+\frac{2}{k-1}\left(\log_{2}\left(\frac{2}{k-1}\right)-\log_{2}\left(\frac{w_{\text{shallow}}}{T}\right)\right)$
Thinking of this as a function of $w_{\text{shallow}}$, the bound is strongest (a local maximum of the right-hand side) when $w_{\text{shallow}}=\frac{2}{(k-1)\log_{2}(T)}$, and weakest at the possible extremes. By our choice of the bins, and size of each bin, we know these extremes are $1/(k-1)\leq w_{\text{shallow}}\leq 3/(k-1)$.
If we assume that $T\geq 3$, the bound is weakest at $w_{\text{shallow}}=3/(k-1)$. This gives
$\hat{H}-H\leq\frac{3\log_{2}(T)}{k-1}+\frac{2}{k-1}\left(\log_{2}\left(\frac{2}{k-1}\right)-\log_{2}\left(\frac{3}{T(k-1)}\right)\right)$
$\hat{H}-H\leq\frac{5\log_{2}(T)}{k-1}-\frac{2\log_{2}(3)}{k-1}$
The second term is small and not very interesting, so we drop it for convenience, completing the proof in the case of $T\geq 3$.
The case of $T<3$ is trivial: our bound only applies when $k\geq 2$, at which point each bin is trivial, so $\hat{H}-H=0$. This completes the proof of Lemma 2.
@@ A.4 Bound for Mixtures
$\hat{H}\leq H+\sum_{j=1}^{s}\frac{5\log_{2}(T)}{(k-2s)}$
$\hat{H}\leq H+\frac{5s\log_{2}(T)}{k-2s}$
@@
Appendix B Derivation of IntMult Bit Cost
@@
$\hat{H}[R_{m}]\approx-p\log_{2}(p)-(m-1)(1-p)\log_{2}(1-p),\qquad p^{3}+(m-1)(1-p)^{3}=\min\left(\frac{\zeta(2)c_{m}}{c},1\right)$
In practice, we use the method of false position to determine this cubic root $p$ in a quick and numericaly stable way. We then plug it in to obtain $\hat{H}[R_{m}]$.

### v2

@@ 3.2.1 Setup
All benchmarks were performed on a bare-metal 96-core Intel Xeon Platinum 8488C CPU. In addition to single-threaded benchmarks, we ran multi-threaded benchmarks to simulate moderate load, using 48 cores such that each active core had its own L2 cache. To minimize measurement noise, we took the median of at least 3 iterations after a warm-up iteration. We limited large datasets to 2 million rows.
We compared Pco’s default configuration versus Blosc2, Parquet, SPDP, Turbo PFor, Vortex [29], and Zstd. We attempted to compare against BtrBlocks directly, but were unable to on account of segmentation faults. Instead we added Vortex, an active open source project whose compression strategy is modeled after BtrBlocks. For Blosc2 we used Shuffle, and for Parquet we considered both Delta and Dictionary. Since Parquet Delta only applies to integers, we always used Dictionary for floats. Blosc2, Parquet, and Turbo PFor were paired with Zstd. We also tested LZ4 and Brotli codecs in place of Zstd, but found that Zstd was strictly better on these datasets.
@@ 48 Zstd 1,178 839 768 893 1,132 912
Table 3: Single-threaded and multi-threaded decompression speeds per thread in MiB/s. Decompression speed varies only slightly between compression levels, so we report its harmonic mean over compression levels. Pco consistently decompressed over 1GiB/s, even in multi-threaded settings.
@@ 3.2.2 Interpretation
Given similar compression time, we found that Pco compressed substantially better than all alternatives on all datasets. Specifically, its default compression level attained 29% to 94% higher compression ratio than other codecs, even when allowing them 50% more compression time (Table 4). In other words, Pco constitutes a 23-48% reduction in storage costs on these datasets. These results matched our intuition that binning would suit numerical data better than LZ codecs do. In the few cases when another codec was able to match the compression ratio of even Pco’s lowest compression level considered (2), Pco compressed faster. In the closest case, Pco at level 2 was 37% faster than Parquet Delta+Zstd at level 1 on the multi-threaded Payments dataset.
@@ 4 Conclusion
Pcodec is a mature compression format that substantially improves state-of-the-art compression on diverse numerical sequences while operating at practical compression and decompression speeds. It accomplishes this via a novel binning procedure, distinct from LZ-based approaches. In addition to leading empirical results, it has strong theoretical results demonstrating rapid convergence to the Shannon entropy of SIID data.
@@ Appendix A Derivation of Binning Bit Cost Bound
Theorem 1 Suppose $X$ is a mixture over domain $\{0,\ldots,T-1\}$ of $s$ disjoint integer distributions, each of which has a monotonic PMF. Then for any $k>2s$, there exists a binning of at most $k$ bins such that the expected binned bit cost $\hat{H}$ of a random draw from $X$ satisfies
$\hat{H}\leq H+\frac{3s\log_{2}(T)}{k-2s}\frac{T}{T-1}$
@@
And less formally, we will show that this binning consists of bins of roughly equal probability. Also informally, the scaling behavior with imbalanced mixture distributions could be improved in subsection A.4.
To sketch the proof, there will be three main steps. First, we will show that on any monotonic distribution, there exists a binning in which every nontrivial bin has probability at most $\frac{3}{2(k-1)}$. Next, we will bound the bit cost when using such a binning. Finally, we will extend this to mixtures of monotonic distributions.
@@ A.1 Definitions
Let $P(x)$ be the PMF of a random integer variable $X$ over a finite, contiguous domain of $T$ distinct $x$ values. Additionally, let a binning with $k$ bins consist of a list of intervals $([a_{1},b_{1}],\ldots,[a_{k},b_{k}])$. Each such bin has various properties:
@@ A.2 Constructive Bound on Bin Weight
If $X$’s PMF is monotonic, then for any $k$, there exists a binning of $X$ with $k$ or fewer bins such that each bin is either trivial or has weight at most $\frac{3}{2(k-1)}$.
We will prove this constructively. Suppose without loss of generality that $X$’s PMF is monotonically decreasing. For any $x$ such that $P(x)>\frac{3}{2(k-1)}$, dedicate a trivial bin to that single $x$ value. Suppose there are $k_{\text{trivial}}$ of these. Following these, all $k_{\text{nontrivial}}$ remaining bins must be nontrivial by monotonicity.
Now we use a simple algorithm to assign nontrivial bins greedily, until their weight would exceed $\frac{3}{2(k-1)}$:
@@
  • For each reamining $x$ value, let $w_{\text{proposed}}=w_{\text{current}}+P(x)$, and
    • If $w_{\text{proposed}}\leq\frac{3}{2(k-1)}$, assign $w_{\text{current}}\leftarrow w_{\text{proposed}}$ and add $x$ to the current bin.
@@
Clearly each such bin has weight of at most $\frac{3}{2(k-1)}$. And for every bin $[a_{i},b_{i}]$ except the last one, we must have that $w_{i}+P(b_{i}+1)>\frac{3}{2(k-1)}$; otherwise we would have included $b_{i}+1$ in the bin. Since the bin is nontrivial, it must contain at least two values, implying that $P(b_{i}-1)+P(b_{i})+P(b_{i}+1)>\frac{3}{2(k-1)}$. By monotonicity, the first two terms must total at least 2/3 of this quantity, giving $w_{i}>\frac{1}{k-1}$.
So we can bound the sum of all trivial and nontrivial bin weights:
$\sum w_{i}>\frac{3k_{\text{trivial}}}{2(k-1)}+\frac{k_{\text{nontrivial}}-1}{k-1}$
$\sum w_{i}>\frac{k_{\text{trivial}}+k_{\text{nontrivial}}-1}{k-1}$
The sum of probabilities must be 1, so
$k-1>k_{\text{trivial}}+k_{\text{nontrivial}}-1$
$k_{\text{trivial}}+k_{\text{nontrivial}}<k$
In other words, we have used at most $k$ bins, each of which is either trivial or weighs at most $\frac{3}{2(k-1)}$. This completes the proof of Lemma 1.
@@ A.3 Bound for Monotonic PMFs
$\hat{H}-H\leq\frac{3\log_{2}(T)}{k-1}\frac{T}{T-1}.$
We will show that the binning from Lemma 1 satisfies this. Without loss of generality, assume $X$’s PMF is monotonically increasing (this is the opposite from the previous section, but makes the indexing simpler). Consider a bin $[a_{i},b_{i}]$. The partial entropy of this bin is
@@
Now we need only consider nontrivial bins. We will further divide these into shallow and deep bins. Let $k_{\text{shallow}}$ be the smallest value such that $\sum_{i=1}^{k_{\text{shallow}}}w_{i}\geq\frac{3}{2(k-1)(T-1)}$. We can rewrite the inequality as
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}&w_{i}\left(\log_{2}(T_{i})-\log_{2}(w_{i})+\log_{2}(q_{i})\right)+\\ \sum_{i=k_{\text{shallow}}+1}^{k_{\text{nontrivial}}}&w_{i}\left(\log_{2}(T_{i})-\log_{2}(w_{i})+\log_{2}(q_{i})\right)\end{aligned}$ (3)
@@
We can use this to bound the deep portion of Equation 3,
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\sum_{i=k_{\text{shallow}}+1}^{k_{\text{nontrivial}}}w_{i}\left(\log_{2}\left(\frac{w_{i}}{q_{i-1}}\right)-\log_{2}(w_{i})+\log_{2}\left(q_{i}\right)\right)$ (4)
@@
Since each of these bins has $w_{i}\leq\frac{3}{2(k-1)}$ by Lemma 1,
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\frac{3}{2(k-1)}\sum_{i=k_{\text{shallow}}+1}^{k_{\text{nontrivial}}}\log_{2}\left(q_{i}\right)-\log_{2}(q_{i-1})$
@@
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\frac{3}{2(k-1)}\left(\log_{2}\left(q_{k_{\text{nontrivial}}}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
Since deep bins are nontrivial, they have $q_{i}\leq\frac{3}{2(k-1)}$, giving
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}\left(\hat{H}_{i}-H_{i}\right)+\frac{3}{2(k-1)}\left(\log_{2}\left(\frac{3}{2(k-1)}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
@@
$\hat{H}-H\leq\sum_{i=1}^{k_{\text{shallow}}}w_{i}\log_{2}(T_{i})+\frac{3}{2(k-1)}\left(\log_{2}\left(\frac{3}{2(k-1)}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
@@
$\hat{H}-H\leq w_{\text{shallow}}\log_{2}(T)+\frac{3}{2(k-1)}\left(\log_{2}\left(\frac{3}{2(k-1)}\right)-\log_{2}\left(q_{k_{\text{shallow}}}\right)\right)$
@@
$\hat{H}-H\leq w_{\text{shallow}}\log_{2}(T)+\frac{3}{2(k-1)}\left(\log_{2}\left(\frac{3}{2(k-1)}\right)-\log_{2}\left(\frac{w_{\text{shallow}}}{T}\right)\right)$
$\hat{H}-H\leq w_{\text{shallow}}\log_{2}(T)+\frac{3}{2(k-1)}\left(\log_{2}\left(\frac{3T}{2(k-1)}\right)-\log_{2}\left(w_{\text{shallow}}\right)\right)$
Thinking of the RHS of this inequality as a function of $w_{\text{shallow}}$, it is clear there is a single local minimum. Therefore this bound must be loosest (a maximum on the RHS) when $w_{\text{shallow}}$ is at an extremal value. Recall that we selected $k_{\text{shallow}}$ to be the smallest value such that $\sum_{i=1}^{k_{\text{shallow}}}w_{i}\geq\frac{3}{2(k-1)(T-1)}$. By this choice, and by the limited size of our nontrivial bins, we know that
$w_{\text{shallow}}\in\left[\frac{3}{2(k-1)(T-1)},\frac{3T}{2(k-1)(T-1)}\right]$
We need only consider these two bounds, $w_{s}=\frac{1}{T-1}$ and $w_{s}=\frac{T}{T-1}$. For convenience, define $w_{s}$ such that $w_{\text{shallow}}=\frac{3}{2(k-1)}w_{s}$. Our inequality becomes
$\hat{H}-H\leq\frac{3}{2(k-1)}\left((w_{s}+1)\log_{2}(T)-\log_{2}\left(w_{s}\right)\right)$
Consider the $w_{s}=w_{s}=\frac{1}{T-1}$ case. We obtain
$\hat{H}-H\leq\frac{3}{2(k-1)}\left(\frac{T}{T-1}\log_{2}(T)+\log_{2}\left(T-1\right)\right)$
Similarly, when $w_{s}=\frac{T}{T-1}$, we get
$\hat{H}-H\leq\frac{3}{2(k-1)}\left(\frac{2T}{T-1}\log_{2}(T)+\log_{2}(T-1)-\log_{2}(T)\right)$
$\hat{H}-H\leq\frac{3}{2(k-1)}\left(\frac{T}{T-1}\log_{2}(T)+\log_{2}(T-1)\right)$
The result is identical in either case, so our bound holds universally. We loosen it slightly to make the result more apparent:
$\hat{H}-H\leq\frac{3\log_{2}(T)}{k-1}\frac{T}{T-1}$
This completes the proof of Lemma 2.
@@ A.4 Bound for Mixtures
$\hat{H}\leq H+\sum_{j=1}^{s}\frac{3\log_{2}(T)}{(k-2s)}\frac{T}{T-1}$
$\hat{H}\leq H+\frac{3s\log_{2}(T)}{k-2s}\frac{T}{T-1}$
@@
Appendix B Derivation of IntMult Bit Cost Approximation
@@
$\hat{H}[R_{m}]\approx-p\log_{2}(p)-(m-1)(1-p)\log_{2}(1-p),\\ \text{s.t.}\quad p^{3}+(m-1)(1-p)^{3}=\min\left(\frac{\zeta(2)c_{m}}{c},1\right)$
In practice, we use the method of false position to determine this cubic root $p$ in a quick and numerically stable way. We then plug it in to obtain $\hat{H}[R_{m}]$.